Questions 4 and 5


media%2F723%2F72345c39-d893-4367-8a31-93 Questions 4 and 5
Answer #1

Utility is maximized when (MU/Price) ratio is same for all goods, where

MU = Change in Total utility (TU) / Change in Quantity (Q)

(Question 4) Option (3)

MUA/PA = 18/3 = 6

MUB/PB = 6/2 = 3

Since (MUA/PA) > (MUB/PB), the consumer should decrease consumption of A and increase consumption of B.

(Question 5) Option (5)

Utility is maximized when (MUX/PX) = (MUY/PY), where

MU = Change in Total utility (TU) / Change in Quantity (Q)

Budget constraint: 44 = 2X + 5Y

QX

TUX

MUX

MUX/PX

QY

TUY

MUY

MUY/PY

1

24

1

85

2

42

18

9

2

130

45

9

3

56

14

7

3

160

30

6

4

66

10

5

4

185

25

5

5

74

8

4

5

200

15

3

6

80

6

3

6

210

10

2

7

84

4

2

7

215

5

1

Therefore, (MUX/PX) = (MUY/PY) is satisfied for following bundles:

(1) (MUX/PX) = (MUY/PY) = 9 when X = 2, Y = 2 and Total cost ($) = 2 x 2 + 5 x 2 = 4 + 10 = 14 < Budget

(2) (MUX/PX) = (MUY/PY) = 5 when X = 4, Y = 4 and Total cost ($) = 2 x 4 + 5 x 4 = 8 + 20 = 28 < Budget

(3) (MUX/PX) = (MUY/PY) = 3 when X = 6, Y = 5 and Total cost ($) = 2 x 6 + 5 x 5 = 12 + 25 = 37 < Budget

(4) (MUX/PX) = (MUY/PY) = 2 when X = 7, Y = 6 and Total cost ($) = 2 x 7 + 5 x 6 = 14 + 30 = 44 = Budget